OpenStudy (anonymous):

The sum of three consecutive integers is 75. Find the integers.

OpenStudy (amistre64):

a+b+c = 75 let a=x-1, b=x, c=x+1

OpenStudy (anonymous):

thank you! how do i solve it though?

OpenStudy (amistre64):

add up a,b,and c in their x forms .... it should look pretty basic after that

OpenStudy (amistre64):

or ... you should know that 25+25+25 = 75 subtract 1 from the start and at it to the end to get 24+25+26 = 75

OpenStudy (anonymous):

thank you so much for your help! :) i was able to solve it and got the answer correct!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

:)